# TabPFN: One Model to Rule Them All?

### v1

@@ 4.2.3 How linear is the TabPFN model?
Table 2: $n=500$, beta-type I, orthogonal design
SNR $s$ TabPFN LASSO
$R^{2}$ Bias2 Var $R^{2}$ Bias2 Var
1.22 1 0.9960 0.0001 0.0017 1.0000 0.0076 0.0020
5 0.9917 0.0007 0.0087 1.0000 0.0201 0.0087
10 0.9774 0.0035 0.0188 1.0000 0.0271 0.0173
20 0.9566 0.0191 0.0502 1.0000 0.0329 0.0362
30 0.9409 0.0435 0.0779 1.0000 0.0309 0.0551
6 1 0.9990 0.0000 0.0004 1.0000 0.0015 0.0004
5 0.9975 0.0001 0.0018 1.0000 0.0041 0.0018
10 0.9939 0.0007 0.0036 1.0000 0.0055 0.0035
20 0.9870 0.0036 0.0091 1.0000 0.0067 0.0074
30 0.9800 0.0089 0.0177 1.0000 0.0062 0.0112
4.3 Robustness-efficiency trade-offs in classification
@@
In our last experiment, we examine the classification performance of TabPFN in the presence of label noise, that is, when the observed labels $\tilde{Y}_{1},\tilde{Y}_{2},\ldots,\tilde{Y}_{n}$ in our training set are corrupted from their true values $Y_{1},Y_{2},\ldots,Y_{n}$, such as by coding errors. For simplicity, we consider homogeneous noise in a binary class setting:
$\mathbb{P}(\widetilde{Y}\neq Y|X=x,Y=y)=\rho,\quad y\in\{0,1\},$
where $\rho$ is the corruption fraction. The goal is to train a classifier to predict the original label. In other words, we want to estimate a mapping $C:\mathbb{R}^{d}\to\{0,1\}$ minimizing the risk
$R(C)=\mathbb{E}[C(X)\neq Y].$
@@
This problem is especially interesting when the conditional distribution of the true label, $p(Y|X=x)$, has a simple parametric form. For example, if $X|Y=y$ is Gaussian, the noise-free classification problem can be solved efficiently via linear discriminant analysis (LDA). On the other hand, the conditional distribution of the corrupted label, $p(\tilde{Y}|X=x)$, does not have the correct functional form, which makes LDA inconsistent. In comparison, Cannings, Fan and Samworth [8] proved that $k$-nearest neighbors ($k$NN) and kernelized support vector machines (SVM) can be robust to label noise. However, these methods, which are based on local averaging, pay a price in terms of being less efficient when the labels are noise free.
@@
We show that TabPFN is able to break this robustness-efficiency trade-off. On the one hand, it remains consistent in the presence of label noise. On the other hand, it is more efficient than $k$NN, SVM and even LDA across all simulation settings, including when there is no label noise.
@@
For each model, we generate an IID sample of size $n+10^{4}$, reserving the first $n$ observations for training and the remaining $10^{4}$ for testing. For the training set, we set $\rho\in\{0.1,\ldots,0.4\}$. The optimal Bayes classifier minimizes this risk:
$C^{\text{Bayes}}(x)=\mathbbm{1}(\eta(x)>0.5),$
where $\eta(x)=\mathbb{P}(Y=1|X=x)$. For any classifier $C^{\prime}$, we define its excess risk as $R(C^{\prime})-R(C^{\text{Bayes}})$.
We compare four classifiers: Bayes, $k$-NN, LDA, and TabPFN. For the Bayes classifier, the true $\eta(x)$ is used in evaluation, and this serves as the gold standard for comparison. Each of the remaining classifiers is trained on two versions of the data: a clean training set and a noise-corrupted version. The resulting classifiers are named accordingly. Details such as hyper-parameter selection are deferred to Appendix F. This process is repeated 1000 times, with Figure 6 showing the average excess risk across repetitions.
@@
Under both generative models, TabPFN (clean) performs as well as the optimal Bayes classifier when the training set size is bigger than $2$K. TabPFN (noisy) shows slightly worse performance than its clean-data counterpart but demonstrates exceptional robustness to label noise in all cases while maintaining high accuracy. Even with noise levels as high as $\rho=0.3$, TabPFN (noisy) nearly matches the optimal classifier given a sufficiently large training set ($\geq 5$K samples). LDA exhibits no robustness under M1 and performs poorly under M2 due to the nonlinear Bayes decision boundary. While the $k$-NN classifier shows some consistency—particularly under M2 with low noise levels—it is less robust than TabPFN in other scenarios. Moreover, Figure 7 demonstrates that TabPFN remains robust even at $\rho=0.4$. Interestingly, we notice that there is a significant performance gap between $\rho=0.3$ and $\rho=0.4$ for small training set sizes.
Figure 7: Excess risk of TabPFN (noisy) under varying $\rho$ values for model M1 (left) and M2 (right).
@@ 5.1 Foundation models
In 2021, Bommasani et al. [6] coined the term “foundation model” to refer to “any model that is trained on broad data (generally using self-supervision at scale) that can be adapted (e.g., fine-tuned) to a wide range of downstream tasks”. They also noted a paradigm shift in artificial intelligence which de-prioritized the development of task-specific models in favor of building these foundation models, with LLMs being the most prominent examples. Until recently, the modality of tabular data has resisted this trend, but the tide may be turning. Hollmann et al. [20] provided a proof-of-concept that, in addition to regression and classification, TabPFN can support “data generation, density estimation, learning reusable embeddings and fine-tuning”, consequently claiming that TabPFN is a tabular foundation model. [^8] These capabilities were not fully investigated by Hollmann et al. [20] and some are not supported by their publicly available code at the time of writing. Our investigations support this claim by showing that simple applications of TabPFN can surpass state-of-the-art approaches to some statistical estimation problems.
These studies are of course preliminary—our numerical experiments made exclusive use of synthetic data—and we do not suggest that TabPFN is a perfect model—it seems to struggle with dense regression, cannot handle $\geq 500$ features, and has MSE curves which seem to plateau for $n\geq 10K$ samples [27]. On the other hand, the evidence suggests that tabular foundation models are not only possible, but can surpass existing modeling approaches on a wide range of statistical tasks, possibly touching on all areas of statistics. In addition, Van Breugel and Van Der Schaar [39] have a bold vision of tabular foundational models being able to perform automatic meta-analyses, out-of-domain simulation, and other tasks for which adequate methodology does not yet exist. This line of thinking—that there may one day be one model to rule them all—is of course contrary to mainstream statistical thinking, which has traditionally emphasized bespoke models for different problems. Just as AI scientists continue to wrestle with the academic and societal consequences of the emergence of foundation models, statisticians may also be forced to confront profound changes in our field.
@@ 5.2 In-context learning
Another way in which TabPFN challenges current statistical thinking is via its illustration of the power of “in-context learning”, whose historical development we now discuss. In a landmark paper, Brown et al. [7] studied the LLM GPT-3 and discovered that it could learn how to perform tasks (e.g., language translation) that it was not explicitly trained to do and without changing any of the model’s parameters, simply by providing examples of the task in a prompt (text input) to the model. They called this ability “in-context learning” (ICL). Garg et al. [15] later framed ICL as a statistical learning problem. Under their framework, a prompt $P$ is a random sequence $(X_{1},f(X_{1}),\ldots,X_{n},f(X_{n}),X_{n+1})$, where $X_{1},\ldots,X_{n+1}$ are drawn IID from a covariate distribution $D_{X}$ and $f$ is drawn from a distribution $D_{f}$ over a function class ${\mathcal{F}}$. A model $M$ is a fixed mapping from prompts to the (real-valued) label space and is said to have learned ${\mathcal{F}}$ in-context (with respect to $(D_{X},D_{f})$) if ${\mathbb{E}}_{P}[l(M(P),f(X_{n+1}))]\leq\epsilon$ for some tolerance $\epsilon$ and loss function $l$. Garg et al. [15] was able to show via numerical experiments that a transformer, trained from scratch on IID prompts generated as described above, could successfully learn simple function classes (e.g. linear functions) in-context, inspiring a wave of theoretical interest in the problem [44, 4, 23].
On the other hand, it seems that Brown et al. [7] and Garg et al. [15]’s definitions of ICL are respectively too broad and too narrow to capture the scope of the learning strategy employed by TabPFN and other recent work claiming to implement ICL to solve statistical estimation problems. These include causal discovery [12], posterior inference for generalized linear models, factor models, and Gaussian mixture models [34], RNA folding time density estimation [36], and Poisson mean estimation under empirical Bayes [38]. None of these works provide a definition of ICL, but we find that the common denominator is that they train a transformer model using a large number of synthetic datasets to learn a mapping on the space of datasets. The target of the mapping can be a causal graph, a posterior distribution on parameter vector, or a Poisson mean vector estimate. In each of these problems, the ICL approach is shown to achieve higher accuracy than state-of-the-art methods, such as variational inference for posterior approximation, or non-parametric maximum-likelihood estimation in the case of empirical Bayes.
@@ 5.3 Future directions
  • Scalability & limitations: Current implementations struggle with dense/high-dimensional regression and $<10$K samples. Addressing this bottleneck is essential for broader applicability. Recent work like [33] offers preliminary solutions, but fundamental constraints remain.
{acks}
[Acknowledgments] This research was enabled in part by support provided by Cedar (cedar.alliancecan.ca) and the Digital Research Alliance of Canada (alliancecan.ca). The authors would like to thank Krishnakumar Balasubramanian, Arik Reuter, Chandan Singh, and Yuqian Zhang for useful discussions.
{funding}
Appendix A Experiment details for semi-supervised estimation
@@ A.1 Simulation settings
    where $X\sim 0.5{\mathcal{N}}_{p}(1,\Sigma)+0.5{\mathcal{N}}_{p}(-1,\Sigma),$ and $\Sigma$ is a matrix with $(i,j)$-th entry equal to $0.5^{|i-j|}$. We set $\alpha_{0}=11$, $\alpha_{1}$ as a vector of all $1$s, and $\alpha_{2}$ as a vector of all $-1$s.
@@ A.2 Additional simulation results
Appendix B Experiment details for causal inference
@@
$\mathbf{A}$ $\min\{0.8,~{}\max\{\sin(\pi X_{1}X_{2}),0.2\}\}$ $\sin(\pi X_{1}X_{2})+2(X_{3}-.5)^{2}+X_{4}+.5X_{5}$$2.2+(X_{1}+X_{2})/$ 2
$\mathbf{B}$5.5.5.5 $\max\{0,X_{1}+X_{2},X_{3}\}+\max\{0,X_{4}+X_{5}\}$ $X_{1}+\log[1+\exp(X_{2})]$
$\mathbf{C}$ $[1+\exp(X_{2}+X_{3})]^{-1}$ $2\log[1+\exp(X_{1}+X_{2}+X_{3})]$ $1$
$\mathbf{D}$ $[1+\exp(-X_{1})+\exp(-X_{2})]^{-1}$$5\max\{0,X_{1}+X_{2}+X_{3}\}+.5\max\{0,X_{4}+X_{5}\}.$ $\max\{0,X_{1}+X_{2}+X_{3}\}-\max\{0,X_{4}+X_{5}\}$
$\mathbf{E}$ $[1+\exp(3X_{2}+3X_{3})]^{-1}$ $5\max\{0,X_{1}+X_{2}\}$ $2\max\{(X_{1}>.1),(X_{2}>.1)\}-$1 ) } - 1
$\mathbf{F}$ $[1+\exp(3X_{2}+3X_{3})]^{-1}$ $5\max\{0,X_{1}+X_{2}\}$ $X_{1}+\log[1+\exp(X_{2})]$
@@
Our experiments are implemented using the econml [5] package in Python, we use the FLAML package [42] to implement all AutoML-based learners. We use the 5-fold cross-validation to choose the optimal algorithms and its corresponding hyperparameters for AutoML-based learners. We randomly generate data from each setting for $100$ times and the average test MSE over $100$ repetitions for CATE is reported in Table 4.
Table 4: The median of test MSE of various CATE estimators over $100$ repetitions. The smallest test MSE is highlighted in boldface and the second smallest test MSE is highlighted with underline.
@@
Appendix C Covariate shift

### v2

@@ 4.1.2 Prediction performance comparison
Table 5: $n=500$, beta-type I, orthogonal design (simple linear regression)
TabPFN LASSO SCAD MCP
SNR $s$ Relevant Irrelevant Relevant Irrelevant Relevant Irrelevant Relevant Irrelevant
Bias2 Var Bias2 Var Bias2 Var Bias2 Var Bias2 Var Bias2 Var Bias2 Var Bias2 Var
1.22 1 0.0000 0.0012 0.0000 0.0000 0.0058 0.0016 0.0000 0.0000 0.0000 0.0012 0.0000 0.0000 0.0000 0.0012 0.0000 0.0000
5 0.0004 0.0081 0.0000 0.0004 0.0172 0.0085 0.0000 0.0006 0.0001 0.0077 0.0000 0.0001 0.0001 0.0077 0.0000 0.0002
10 0.0028 0.0196 0.0000 0.0025 0.0244 0.0174 0.0000 0.0020 0.0002 0.0180 0.0000 0.0007 0.0002 0.0179 0.0000 0.0007
20 0.0175 0.0500 0.0004 0.0099 0.0303 0.0360 0.0001 0.0079 0.0005 0.0400 0.0001 0.0051 0.0005 0.0407 0.0001 0.0049
30 0.0387 0.0783 0.0018 0.0177 0.0301 0.0579 0.0002 0.0166 0.0011 0.0707 0.0002 0.0146 0.0011 0.0719 0.0002 0.0142
6 1 0.0000 0.0003 0.0000 0.0000 0.0022 0.0002 0.0000 0.0000 0.0000 0.0002 0.0000 0.0000 0.0000 0.0002 0.0000 0.0000
5 0.0000 0.0016 0.0000 0.0001 0.0038 0.0017 0.0000 0.0001 0.0000 0.0015 0.0000 0.0000 0.0000 0.0016 0.0000 0.0000
10 0.0004 0.0037 0.0000 0.0006 0.0053 0.0035 0.0000 0.0004 0.0000 0.0033 0.0000 0.0001 0.0000 0.0033 0.0000 0.0001
20 0.0022 0.0089 0.0001 0.0028 0.0070 0.0073 0.0000 0.0014 0.0001 0.0067 0.0000 0.0002 0.0001 0.0067 0.0000 0.0002
30 0.0055 0.0168 0.0003 0.0061 0.0079 0.0116 0.0000 0.0027 0.0001 0.0110 0.0000 0.0004 0.0001 0.0110 0.0000 0.0004
4.1.3 Descriptive accuracy
Beyond predictive accuracy, we also consider the descriptive accuracy of the TabPFN model–how well does it reflect the sparse linear generating model and how it may be used to estimate the true regression coefficients. We can visually inspect the fitted TabPFN model using Accumulated Local Effects (ALE) plots, which summarize the dependence of the model on each individual feature [2]. [^9] ALE plots avoid both the confounding associated with conditional dependence plots and the out-of-sample instability associated with marginal dependence plots. Figure 21 in the Appendix exhibits the results when the features are orthogonal. We observe that for moderate to high SNR ($\geq 1.22$), almost all ALE plots of relevant features are roughly linear. In addition, when $n=500$, the ALE plots for the irrelevant features display much less variation, which suggests that the TabPFN model is indeed well-approximated by a low-dimensional linear surface. For smaller SNR, the ALE plots are generally nonlinear and are not shown.
The linearity of the ALE plots suggests a two-step pipeline for using TabPFN to estimate the true regression coefficients: We first fit a TabPFN model to the data and then fit an ordinary least squares model to the TabPFN fitted values. In Table 5, we compare the accuracy of this procedure to that of LASSO, SCAD, and MCP, further breaking down the error into squared bias and variance. For smaller values of $s$, we observe that the TabPFN performs similarly to SCAD and MCP: Its coefficients have much smaller bias compared to LASSO while having comparable variance, thus avoiding the well-known tendency of LASSO to bias its estimates via shrinkage.
4.2 Adaptivity in classification
@@
We examine the classification performance of TabPFN in the presence of label noise, that is, when the observed labels $\tilde{Y}_{1},\tilde{Y}_{2},\ldots,\tilde{Y}_{n}$ in our training set are corrupted from their true values $Y_{1},Y_{2},\ldots,Y_{n}$, such as by coding errors. For simplicity, we consider homogeneous noise in a binary class setting: $\mathbb{P}(\widetilde{Y}\neq Y|X=x,Y=y)=\rho,y\in\{0,1\},$ where $\rho$ is the corruption fraction. The goal is to train a classifier to predict the original label, namely, we want to estimate a mapping $C:\mathbb{R}^{d}\to\{0,1\}$ minimizing the risk $R(C)=\mathbb{E}[C(X)\neq Y].$
@@
This problem is especially interesting when the conditional distribution of the true label, $p(Y|X=x)$, has a simple parametric form. For example, if $X|Y=y$ is Gaussian, the noise-free classification problem can be solved efficiently via linear discriminant analysis (LDA). On the other hand, the conditional distribution of the corrupted label, $p(\tilde{Y}|X=x)$, does not have the correct functional form, which makes LDA inconsistent. In comparison, Cannings, Fan and Samworth [10] proved that $k$-nearest neighbors (kkNN) and kernelized support vector machines (SVM) can be robust to label noise. However, these methods, which are based on local averaging, pay a price in terms of being less efficient when the labels are noise free.
@@
We show that TabPFN is able to break this robustness-efficiency trade-off. On one hand, it remains consistent in the presence of label noise. On the other hand, it is more efficient than kkNN, SVM and even LDA across all simulation settings, including noiseless settings.
@@
For each model, we generate an IID sample of size $n+10^{4}$, reserving the first $n$ observations for training and the remaining $10^{4}$ for testing. For the training set, we set $\rho\in\{0.1,\ldots,0.4\}$. The optimal Bayes classifier minimizes this risk: $C^{\text{Bayes}}(x)=\mathbbm{1}(\eta(x)>0.5),$ where $\eta(x)=\mathbb{P}(Y=1|X=x)$. For any classifier $C^{\prime}$, we define its excess risk as $R(C^{\prime})-R(C^{\text{Bayes}})$.
We compare three classifiers: Bayes, $k$-NN, LDA, and TabPFN. For each training dataset, each classifier trained on two versions of the data: a clean training set and a noise-corrupted version, hence yielding a total of six classification models to be evaluated. Details such as hyper-parameter selection are deferred to Appendix F. This process is repeated 1000 times, with Figure 6 showing the average excess risk across repetitions.
@@
Under both generative models, TabPFN (clean) performs as well as the optimal Bayes classifier when the training set size is bigger than 22K. TabPFN (noisy) shows slightly worse performance than its clean-data counterpart but demonstrates exceptional robustness to label noise in all cases while maintaining high accuracy. Even with noise levels as high as $\rho=0.3$, TabPFN (noisy) nearly matches the optimal classifier given a sufficiently large training set ($\geq 5$K samples). LDA exhibits no robustness under M1 and performs poorly under M2 due to the nonlinear Bayes decision boundary. While the $k$-NN classifier shows some consistency—particularly under M2 with low noise levels—it is less robust than TabPFN in other scenarios. Moreover, Figure 7 demonstrates that TabPFN remains robust even at $\rho=0.4$. Interestingly, we notice that there is a significant performance gap between $\rho=0.3$ and $\rho=0.4$ for small training set sizes.
Figure 7: TabPFN (noisy) excess risk for varying $\rho$ values under M1 (left) and M2 (right).
@@ 5.1 Foundation models
In 2021, Bommasani et al. [7] coined the term “foundation model” to refer to “any model that is trained on broad data (generally using self-supervision at scale) that can be adapted (e.g., fine-tuned) to a wide range of downstream tasks”. They also noted a paradigm shift in artificial intelligence which de-prioritized the development of task-specific models in favor of building these foundation models, with LLMs being the most prominent examples. Until recently, the modality of tabular data has resisted this trend, but the tide may be turning. Hollmann et al. [30] provided a proof-of-concept that, in addition to regression and classification, TabPFN can support “data generation, density estimation, learning reusable embeddings and fine-tuning”, consequently claiming that TabPFN is a tabular foundation model. [^10] These capabilities were not fully investigated by Hollmann et al. [30] and some are not supported by their publicly available code at the time of writing. Our investigations support this claim by showing that simple applications of TabPFN can surpass state-of-the-art approaches to some statistical problems.
These studies are of course preliminary and we do not suggest that TabPFN is a perfect model—it seems to struggle with dense regression, cannot handle $\geq 500$ features, and has MSE curves which seem to plateau for $n\geq 10K$ samples [38]. On the other hand, the evidence suggests that tabular foundation models are not only possible, but have the potential to reshape how a wide range of statistical tasks are approached. In addition, Van Breugel and Van Der Schaar [52] have a bold vision of tabular foundational models being able to perform automatic meta-analyses, out-of-domain simulation, and other tasks for which adequate methodology does not yet exist. This line of thinking—that there may one day be one model to rule them all—is of course contrary to mainstream statistical thinking, which has traditionally emphasized bespoke models for different problems. Just as AI scientists wrestle with the academic and societal consequences of foundation models, statisticians may also be forced to confront profound changes in our field.
@@ 5.2 In-context learning
Another way in which TabPFN challenges current statistical thinking is via its illustration of the power of “in-context learning”, whose historical development we now discuss. In a landmark paper, Brown et al. [9] studied the LLM GPT-3 and discovered that it could learn how to perform tasks (e.g., language translation) that it was not explicitly trained to do and without changing any of the model’s parameters, simply by providing examples of the task in a prompt (text input) to the model. They called this ability “in-context learning” (ICL). Garg et al. [22] later framed ICL as a statistical learning problem. Under their framework, a prompt $P$ is a random sequence $(X_{1},f(X_{1}),\ldots,X_{n},f(X_{n}),X_{n+1})$, where $X_{1},\ldots,X_{n+1}$ are drawn IID from a covariate distribution $D_{X}$ and $f$ is drawn from a distribution $D_{f}$ over a function class $\mathcal{F}$}. A model $M$ is a fixed mapping from prompts to the (real-valued) label space and is said to have learned $\mathcal{F}$} in-context (with respect to $(D_{X},D_{f})$) if ${\mathbb{E}}_{P}[l(M(P),f(X_{n+1}))]\leq\epsilon$ for some tolerance $\epsilon$ and loss function $l$. Garg et al. [22] was able to show via numerical experiments that a transformer, trained from scratch on IID prompts generated as described above, could successfully learn simple function classes (e.g. linear functions) in-context, inspiring a wave of theoretical interest in the problem (see for instance Zhang, Frei and Bartlett [57], Bai et al. [5], Kim, Nakamaki and Suzuki [33]).
On the other hand, it seems that Brown et al. [9] and Garg et al. [22]’s definitions of ICL are respectively too broad and too narrow to capture the scope of the learning strategy employed by TabPFN and other recent work claiming to implement ICL to solve statistical estimation problems. These include causal discovery [15], posterior inference for generalized linear models, factor models, and Gaussian mixture models [45], RNA folding time density estimation [48], and Poisson mean estimation under empirical Bayes [50]. None of these works provide a definition of ICL, but we find that the common denominator is that they train a transformer model using a large number of synthetic datasets to learn a mapping on the space of datasets. The target of the mapping can be a causal graph, a posterior distribution on parameter vector, or a Poisson mean vector estimate. In each problem, the ICL approach achieves higher accuracy than SOTA methods, such as variational inference for posterior approximation, or non-parametric maximum-likelihood estimation in the case of empirical Bayes.
@@ 5.3 Future directions
  • Scalability & limitations: Current implementations struggle with dense/high-dimensional regression and >10>10K samples. Addressing this bottleneck is essential for broader applicability. Recent work  [44, 25] offers preliminary solutions, but fundamental constraints remain.
Appendix A Experiment details for study I
@@ A.1 Simulation settings
    where $X\sim 0.5{\mathcal{N}}_{p}(1,\Sigma)+0.5{\mathcal{N}}_{p}(-1,\Sigma),$ and $\Sigma$ is a matrix with $(i,j)$-th entry equal to $0.5^{|i-j|}$. We set $\alpha_{0}=11$, $\alpha_{1}$ as a vector of all $1$s, and $\alpha_{2}$ as a vector of all 1-1s.
@@ A.2 Additional simulation results
A.3 Real data comparison
The datasets are obtained from OpenML. We used the dataset search interface at www.openml.org/, selecting datasets with fewer than 10 features and more than 1,000 instances, and verified that all features are numerical. We then downloaded datasets from the top of this filtered list. The restriction to fewer than $10$ features is imposed because the SLZ implementation can exhibit numerical instability when the dimensionality is large. The requirement of a sufficiently large number of instances ensures that the oracle estimator provides a reasonable approximation to the true parameter values. To maintain computational efficiency, we further restricted datasets to have fewer than 50,000 instances. Even after this filtering, we occasionally encountered matrix-inversion issues in the SLZ implementation; when such cases occurred, we skipped the problematic dataset and selected the next one on the list. This leads to the following datasets for regression and classification respectively:
  • Regression: kin8nm (ID:189), house_8L (ID: 218), puma8NH (ID: 225), BNG (stock) (ID: 1200), cmc (ID: 45052)
  • Classification: BNG(breast-w) (ID:251), kin8nm (ID:807), puma8NH (ID: 816), mozilla4 (ID: 1046), wilt (ID: 40983)
Appendix B Experiment details for study II
@@ B.1 Simulation settings and additional results
$\mathbf{A}$ $\min\{0.8,~\max\{\sin(\pi X_{1}X_{2}),0.2\}\}$ $\sin(\pi X_{1}X_{2})+2(X_{3}-.5)^{2}+X_{4}+.5X_{5}$ $2+(X_{1}+X_{2})/2$
$\mathbf{B}$ .5.5 $\max\{0,X_{1}+X_{2},X_{3}\}+\max\{0,X_{4}+X_{5}\}$ $X_{1}+\log[1+\exp(X_{2})]$
$\mathbf{C}$ $1+\exp(X_{2}+X_{3})]^{-1}$ $2\log[1+\exp(X_{1}+X_{2}+X_{3})]$ 11
$\mathbf{D}$ $1+\exp(-X_{1})+\exp(-X_{2})]^{-1}$ $5\max\{0,X_{1}+X_{2}+X_{3}\}+.5\max\{0,X_{4}+X_{5}\}$ $\max\{0,X_{1}+X_{2}+X_{3}\}-\max\{0,X_{4}+X_{5}\}$
$\mathbf{E}$ $1+\exp(3X_{2}+3X_{3})]^{-1}$ $5\max\{0,X_{1}+X_{2}\}$ $2\max\{(X_{1}>.1),(X_{2}>.1)\}-1$
$\mathbf{F}$ $1+\exp(3X_{2}+3X_{3})]^{-1}$ $5\max\{0,X_{1}+X_{2}\}$ $X_{1}+\log[1+\exp(X_{2})]$
@@
Our experiments are implemented using the econml [6] package in Python, we use the FLAML package [55] to implement all AutoML-based learners. We use the 5-fold cross-validation to choose the optimal algorithms and its corresponding hyperparameters for AutoML-based learners. We randomly generate data from each setting for $100$ times and the average test MSE over $100$ repetitions for CATE is reported in Table 7 and Table 8.
Table 7: The median of test MSE of various CATE estimators over $100$ repetitions setup A–C. The smallest test MSE is highlighted in boldface and the second smallest test MSE is highlighted with underline.
@@
Table 8: The median of test MSE of various CATE estimators over $100$ repetitions under setup D–F. The smallest test MSE is highlighted in boldface and the second smallest test MSE is highlighted with underline.
Setup N $\sigma^{2}$ DR S T X R Oracle
@@
B.2 Real data comparison
Data. The Infant Health and Development Program (IHDP) [8] dataset contains $n=4302$ subjects with $8$ covariates capturing demographic, prenatal, and birth-related characteristics, a binary treatment indicating participation in a specialist early-childhood intervention program, and continuous outcomes representing cognitive test scores measured at age $3$. We use a collection of semi-synthetic versions of the dataset that were created for the ACIC 2017 Causal Inference Benchmark challenge described in Hahn, Dorie and Murray [26], where treatment assignments and potential outcomes are simulated conditional on the real-world covariates. This design ensures access to the ground-truth treatment effects, allowing quantitative comparison across estimators via the MSE of the CATE. We accessed the data via Richard Hahn’s personal website: https://math.la.asu.edu/~prhahn/.
Each semi-synthetic dataset is generated according to one of 24 data-generating processes (DGPs) that were chosen to systematically vary structural and noise properties to probe robustness of causal estimators. Specifically, the DGPs are organized into three broad settings for the error structure:
  • IID: Errors are generated independently across individuals, corresponding to the standard assumption of homogeneous, uncorrelated residuals.
  • Non-additive: Treatment effects and outcomes are generated via nonlinear functions and interaction terms, inducing complex, non-additive response surfaces.
  • Group Corr: The error terms include intra-group correlation components to mimic unobserved group-level dependencies (e.g., hospitals or study sites), introducing residual correlation beyond observed covariates and thereby violating the IID assumption.
A fourth error type, heteroskedastic case, was later discovered by Hahn, Dorie and Murray [26] to contain inconsistencies and hence omitted from the benchmark comparison.
Within each setting, eight distinct DGPs are defined to explore combinations of three factors that affect causal inference difficulty: the effect magnitude $\xi$ (large vs. small), the noise level $\eta$ (low vs. high), and the selection strength $\kappa$ (weak vs. strong confounding). Overall, these yield $3$ settings $\times$ 88 scenarios $\times$ $250$ replications =6,000=6{,}000 semi-synthetic datasets derived from the IHDP covariates.
Evaluation. We compare four TabPFN-based meta-learners (S-Learner, T-Learner, X-Learner, and DR-Learner) with the 22 estimators reported in the benchmark challenge. For a fair comparison, we only include methods with available results across all 6,0006{,}000 datasets, resulting in 16 baseline approaches. For all TabPFN-based methods, we use the TabPFN classifier to estimate the propensity score model and the TabPFN regressor to fit the outcome model, as well as any auxiliary models with continuous responses. The evaluation protocol follows that of Hahn, Dorie and Murray [26], and we estimate the RMSE of the CATEs based on 250 repetitions under each scenario. The results for the baseline approaches were taken directly from Richard Hahn’s personal website and were not recalculated.
The comparison of different methods across the 24 scenarios is presented in Figure 14. TabPFN-based methods, especially X+PS and X learners shown strong performance in almost all scenarios.
Figure 14: Performance comparison of TabPFN-based CATE estimators against SOTA methods on the ACIC 2017 benchmark. Results are shown across panels organized by error type, effect magnitude ($\xi$), noise level ($\eta$), and selection strength ($\kappa$). Colored bars represent meta-learners built on TabPFN: S-Learner, T-Learner, X-Learner, DR-Learner, S+PS-Learner, T+PS-Learner, and X+PS-Learner. The BCF is shown for reference as the shaded yellow bar.
Appendix C Experimental details for study III
C.1 Additional simulation results